# Closing the Gap: Achieving Better Accuracy-Robustness Tradeoffs Against Query-Based Attacks

### 15 Dec 2023 v1

@@ Introduction
85 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.52 0.95 0.36 0.94 0.40 0.94 0.67 0.94 0.44 0.92 0.83 0.92 0.65 0.92 0.73 0.59 0.79 0.58 0.47 0.58 0.90 0.57 0.51 0.56 0.94 0.56 0.85 0.56 0.49 0.56 0.85 0.80 0.77 0.80 0.80 0.79 0.79 0.79 0.79 0.79 0.77 0.78 0.78 0.78 0.78 0.78 0.78
75 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.37 0.94 0.57 0.94 0.49 0.94 0.40 0.93 0.61 0.91 0.84 0.90 0.71 0.90 0.67 0.59 0.45 0.58 0.78 0.57 0.81 0.56 0.50 0.56 0.51 0.55 0.60 0.54 0.89 0.54 0.82 0.80 0.82 0.80 0.79 0.79 0.79 0.78 0.76 0.77 0.75 0.76 0.70 0.76 0.72 0.76 0.72
60 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.41 0.94 0.50 0.94 0.31 0.94 0.44 0.93 0.36 0.90 0.77 0.87 0.60 0.87 0.60 0.58 0.65 0.57 0.52 0.56 0.89 0.55 0.60 0.53 0.50 0.52 0.50 0.52 0.55 0.52 0.83 0.80 0.80 0.79 0.79 0.79 0.76 0.78 0.76 0.76 0.74 0.74 0.65 0.74 0.64 0.74 0.64
50 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.55 0.94 0.46 0.94 0.32 0.94 0.38 0.93 0.55 0.88 0.75 0.85 0.60 0.85 0.64 0.58 0.49 0.57 0.82 0.56 0.56 0.54 0.88 0.52 0.52 0.51 0.56 0.50 0.60 0.50 0.83 0.80 0.80 0.79 0.77 0.78 0.75 0.77 0.75 0.76 0.75 0.73 0.68 0.73 0.67 0.73 0.67
Table 1: $\mathsf{RA}$ and $\mathsf{CA}$ based on the defense parameter $\nu$ and the threshold $\tau$. We mark in bold the improvements beyond the Pareto frontier achieved in vanilla constructs (i.e., when $\tau=1.0$). See the complete results in the Appendix.
@@
Let $S\subset\mathcal{X}\times\mathcal{Y}$ denote the set of (labeled) genuine samples provided to the attacker, and let $n:=\lvert S\rvert$, let $Q$ denote the query budget, and let $\varepsilon$ denote the distortion budget. To compute the attack success rate (ASR) in practice. we determine the number of successful adversarial samples generated by the attacker:
@@
where $\mathcal{A}(S)$ denotes the set of candidate adversarial samples output by $\mathcal{A}$ in a run of the attack on input $S$. The ASR is defined as $\mathsf{ASR}:=\nicefrac{{n_{\mathrm{succ}}}}{{n}}$, i.e., the ratio of successful adversarial samples. The complement of the $\mathsf{ASR}$ is the robust accuracy ($\mathsf{RA}$) of the classifier, $\mathsf{RA}=1-\mathsf{ASR}$.
@@
To evaluate the effectiveness of our approach, we instantiate our proposal with three existing defenses $\mathsf{D}$, namely RND, RCR, and JPEG compression, and we empirically measure the accuracy and robust accuracy of the resulting classifiers $F_{\mathsf{D},\tau}$ against state-of-the-art decision-based attacks PSJA and SurFree in the untargeted setting.
@@
Throughout our evaluation, we evaluate PSJA against randomized defenses, i.e., RND, RCR, and SurFree, against the deterministic defense, i.e., JPEG. We always measure the clean accuracy (or $\mathsf{CA}$) of the classifier over the entire test set of each respective dataset and the robust accuracy ($\mathsf{RA}$) on $n=100$ randomly selected images from the test set that are correctly classified by the undefended classifier.
Our experiments have been conducted on a server equipped with two AMD EPYC 7542 CPUs, two A40 GPUs, and 256GB RAM. The system runs Ubuntu 22.04., Python 3.9, PyTorch 1.13.0, and CUDA 11.7.
@@
(d) PSJA, CIFAR-100, $\mathsf{D}$ = RND
(e) PSJA, CIFAR-100, $\mathsf{D}$ = RCR
(f) SurFree, CIFAR-100, $\mathsf{D}$ = JPEG
Figure 3: Pareto frontier when applying our approach in conjunction with RND, RCR, and JPEG on CIFAR-10 and CIFAR-100 using the attacks PSJA and SurFree.
Figure 4: Pareto frontier when applying our approach in conjunction with JPEG on ImageNet using SurFree.
@@
In Figure 3, we use the Pareto frontier to showcase the trade-off improvements of our proposal compared to the baseline relying only on the defense parameter $\nu$. We plot the points $(\mathsf{CA}(\omega^{*}),\mathsf{RA}(\omega^{*}))$ for empirically-determined Pareto-optimal solutions $\omega^{*}$ for the baseline defense ($\omega^{*}=\nu^{*}$, solid line) and for our approach ($\omega^{*}=(\nu^{*},\tau^{*})$, dotted line). The accuracy-robustness tradeoffs achieved by our proposal for selected threshold values are highlighted in grey. The specific combinations of $\tau,\nu$ that outperform the baseline tradeoffs (${\tau=1.0}$) are highlighted in Table 1. Note that for RCR the crop size for CIFAR-10/100 and ImageNet are shown due to the differently sized input dimensions.
@@
When evaluated against PSJA on the CIFAR-10 dataset, our proposal combined with the RND defense can provide an improvement in $\mathsf{RA}$ while preserving the  $\mathsf{CA}$ of the baseline. For example, it improves $\mathsf{RA}$ by $8\%$ (out of a maximum of $9\%$) for $\mathsf{CA}=0.95$, and it improves $\mathsf{RA}$ by up to $9\%$ for a trade-off of at most $1\%$ decrease in $\mathsf{CA}$. Compared to the baseline ($\tau=1.0,\nu=0.07$), we improve $\mathsf{CA}$ by $29\%$, while decreasing $\mathsf{RA}$ by just $1\%$ as seen in Figure 2(a).
@@
Our results on CIFAR-100 shown in Figures 2(d) and 2(e) are consistent with our CIFAR-10 results. More precisely, we measure an improvement of $\mathsf{RA}$ of up to $7\%$ for a $\mathsf{CA}$ of $58\%$ when evaluating with the RND defense. Compared to the baseline with a $\mathsf{CA}/\mathsf{RA}$ of $37\%$, our approach improves the $\mathsf{CA}$ by $16\%$ and $\mathsf{RA}$ by $1\%$. When considering the RCR defense in Figure 2(e), our approach can improve $\mathsf{CA}$ by $19\%$ for the same $\mathsf{RA}$ of $43\%$. When $\mathsf{CA}=0.53$, our approach yields an improvement in $\mathsf{RA}$ by $8\%$. We discuss the ImageNet results in the Appendix.
@@
Figure 2(c), Figure 2(f), and Figure 4 depict the evaluation results against the SurFree attack, for which our proposal can obtain consistent improvements compared to the baseline. At no decrease in $\mathsf{CA}$, we observe significant improvements in $\mathsf{RA}$. For CIFAR-10, it increases by $34\%$ and for CIFAR-100 by $21\%$. This trend is also observable for ImageNet – we obtain an increase of $\mathsf{RA}$ of $21\%$ with a negligible decrease of $\mathsf{CA}$ by $1\%$. When fixing $\mathsf{RA}$ to $67\%$, we obtain an improvement in $\mathsf{CA}$ of $4\%$ for CIFAR-10. Moreover, we see an improvement of $\mathsf{CA}$ by $3\%$ at a fixed $\mathsf{RA}=0.82$ for CIFAR-100.
@@
We now compare our proposal to various training-time defenses, which require the costly retraining of a model and cannot be retrofitted to existing models. We compare against OAT (Wang et al. 2020) because it is the only reconfigurable, i.e., adjustment of accuracy-robustness tradeoff after training, adversarial-training solution with an open-source implementation. In addition, we include traditional, non-adjustable, training-time defenses, such as PNI (He, Rakin, and Fan 2019), RSE (Liu et al. 2018), and a state-of-the-art robust model (AT) (Gowal et al. 2020) from RobustBench for further comparison. We also consider the previously considered vanilla test-time defenses RND, RCR, and JPEG with $\tau=1.0$, i.e., traditional deployment of defenses for this comparison.
Training-time Test-time
Baseline
@@
Table 2: $\mathsf{RA}$ and $\mathsf{CA}$ for training- and test-time defenses on CIFAR-10.
In Table 2, we evaluate $\mathsf{RA}$ at those values of $\mathsf{CA}$ obtained with all defenses against PSJA and SurFree on CIFAR-10. We do not report results when a defense does not yield those values of $\mathsf{CA}$. In comparison with traditionally obtainable tradeoffs, our results show that our approach consistently improves $\mathsf{RA}$ for all $\mathsf{CA}$ by up to $38\%$ with a minimum increase of $9\%$. On average, we obtain an improvement of $\sim 20\%$ in $\mathsf{RA}$ for PSJA and $\sim 11\%$ for SurFree.
Against SurFree, our proposal manages to outperform all training-/test-time based defenses, outperforming AT by $8\%$, while being completely training-free. For $\mathsf{CA}=0.88$, we observe a difference to the $\mathsf{RA}$ of OAT of just $7\%$. For ${\mathsf{CA}=0.84}$, a larger difference of $23\%$ in $\mathsf{RA}$ to the one of PNI can be measured. We conclude that, although it is training-free, our proposal manages to outperform or closely perform similarly to existing defenses against SurFree.
In the case of PSJA, we observe that training-time defenses outperform our approach by $\sim 30\%$ in $\mathsf{RA}$. Compared to OAT, this difference is reduced to just $18\%$. Our approach, however, results in significant improvements compared to all training-free defenses.
@@
Our proposal relies on two main parameters: the confidence threshold $\tau$ and the genuine defense’s parameter $\nu$. To evaluate the impact of each of these parameters on robustness, we present an ablation study. We detail our results in Table 1, where we show the achieved $\mathsf{CA}$ (top of each row) and $\mathsf{RA}$ (bottom of each row) for different values of $\tau$ and $\nu$. We highlight the values that result in an improvement of the Pareto frontier in bold. When $\tau=1.0$, our proposal instantiates the baseline since the defense is always enabled.
@@
We vary the value of $\tau\in\{0.0,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.97,0.99,1.0\}$. For the CIFAR-10 dataset, we first consider the RND defense with noise $\nu=0.05$, which offers the highest $\mathsf{RA}=0.71$ at $\mathsf{CA}=0.82$ and the best tradeoffs. As we reduce the value of $\tau$, we observe a significant improvement in $\mathsf{CA}$, increasing from $71\%$ when $\tau\geq 0.99$ to $87\%$ when $\tau=0.97$. By further reducing $\tau$ to $0.8$, our proposal achieves $94\%$ $\mathsf{CA}$, fairly close to the original $\mathsf{CA}$ of $95\%$. On the other hand, $\mathsf{RA}$ starts at $71\%$ for the baseline, it remains consistent at $94\%$ when $\tau=0.99$, and decreases to $64\%$ when $\tau=0.97$. For $\tau\leq 0.8$, $\mathsf{RA}$ remains at $59\%$ and then drops sharply to a $\mathsf{RA}$ as low as $43\%$. The Pareto frontier for the fixed $\nu$ and varying $\tau$ can be found in Figure 4(a).
For RCR, we show the Pareto frontier for $\nu=22$ in Figure 4(c) and for JPEG for $\nu=85$ in Figure 4(e). For both defenses, we observe consistent improvements in the baseline tradeoffs. We argue that such a noise level introduced by these two defenses is similar to $\nu=0.05$ for the RND defense. These observations are not exclusive to CIFAR-10 but can also be clearly seen with CIFAR-100 with $\nu=26$ for RCR, while it is observable for $\nu=0.05$ and $\nu=85$ for RND and JPEG, respectively.
@@
(d) PSJA, $\mathsf{D}$ = RCR,
$\tau=0.7$
(e) SurFree, $\mathsf{D}$ = JPEG,
$\nu=85$
@@
Figure 5: Pareto frontier for threshold $\tau$ and noise level $\nu$ for CIFAR-10.

### 21 Mar 2024 v2

@@ Introduction
85 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.52 0.95 0.36 0.94 0.40 0.94 0.67 0.92 0.83 0.92 0.65 0.92 0.73 0.59 0.79 0.58 0.47 0.58 0.90 0.57 0.51 0.56 0.94 0.56 0.85 0.56 0.49 0.56 0.85
75 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.37 0.94 0.57 0.94 0.49 0.94 0.40 0.91 0.84 0.90 0.71 0.90 0.67 0.59 0.45 0.58 0.78 0.57 0.81 0.56 0.50 0.56 0.51 0.55 0.60 0.54 0.89 0.54 0.82
60 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.41 0.94 0.50 0.94 0.31 0.94 0.44 0.90 0.77 0.87 0.60 0.87 0.60 0.58 0.65 0.57 0.52 0.56 0.89 0.55 0.60 0.53 0.50 0.52 0.50 0.52 0.55 0.52 0.83
50 $\mathsf{CA}$ $\mathsf{RA}$ 0.95 0.55 0.94 0.46 0.94 0.32 0.94 0.38 0.88 0.75 0.85 0.60 0.85 0.64 0.58 0.49 0.57 0.82 0.56 0.56 0.54 0.88 0.52 0.52 0.51 0.56 0.50 0.60 0.50 0.83
Table 1: Selected results for $\mathsf{RA}$ and $\mathsf{CA}$ based on the defense parameter $\nu$ and the threshold $\tau$ in the CIFAR-10 and the CIFAR-100 datasets. Improvements beyond the Pareto frontier achieved in vanilla constructs (i.e., when $\tau=1.0$) are bolded.
Dataset ImageNet
$\nu$ / $\tau$ 0.0 0.3 0.5 0.6 0.9 1.0
PSJA with RND
0.01 $\mathsf{CA}$ $\mathsf{RA}$ 0.81 0.97 0.81 0.97 0.80 0.98 0.80 0.97 0.80 0.98 0.80 0.97
0.02 $\mathsf{CA}$ $\mathsf{RA}$ 0.81 0.97 0.81 0.97 0.80 0.98 0.80 0.98 0.80 0.97 0.80 0.97
0.05 $\mathsf{CA}$ $\mathsf{RA}$ 0.81 0.97 0.80 0.96 0.79 0.94 0.78 0.94 0.77 0.91 0.76 0.91
SurFree with JPEG
85 $\mathsf{CA}$ $\mathsf{RA}$ 0.81 0.61 0.80 0.77 0.79 0.79 0.79 0.79 0.78 0.78 0.78 0.78
75 $\mathsf{CA}$ $\mathsf{RA}$ 0.81 0.61 0.80 0.82 0.79 0.79 0.78 0.76 0.76 0.70 0.76 0.72
60 $\mathsf{CA}$ $\mathsf{RA}$ 0.81 0.61 0.80 0.80 0.79 0.76 0.78 0.76 0.74 0.65 0.74 0.64
Table 2: Selected results for $\mathsf{RA}$ and $\mathsf{CA}$ based on the defense parameter $\nu$ and the threshold $\tau$ in the ImageNet dataset. Our complete results are included in the Appendix.
@@
Let $S\subset\mathcal{X}\times\mathcal{Y}$ denote the set of (labeled) genuine samples provided to the attacker, let $n:=\lvert S\rvert$, let $Q$ denote the query budget, and let $\varepsilon$ denote the distortion budget. To compute the attack success rate ($\mathsf{ASR}$) in practice. we determine the number of successful adversarial samples generated by the attacker:
@@
where $\mathcal{A}(S)$ denotes the set of candidate adversarial samples output by $\mathcal{A}$ in a run of the attack on input $S$. The $\mathsf{ASR}$ is defined as $\mathsf{ASR}:=\nicefrac{{n_{\mathrm{succ}}}}{{n}}$, i.e., the ratio of successful adversarial samples. The complement of the $\mathsf{ASR}$ is the robust accuracy ($\mathsf{RA}$) of the classifier, i.e., $\mathsf{RA}=1-\mathsf{ASR}$.
@@
To evaluate the effectiveness of our approach, we instantiate our proposal with three existing defenses $\mathsf{D}$, namely RND, RCR, and JPEG compression, and empirically measure the accuracy and robust accuracy of the resulting classifiers $F_{\mathsf{D},\tau}$ against state-of-the-art decision-based attacks PSJA and SurFree in the untargeted setting.
@@
Throughout our evaluation, we evaluate PSJA against randomized defenses, i.e., RND, RCR, and SurFree against the deterministic defense, i.e., JPEG. We always measure the clean accuracy (or $\mathsf{CA}$) of the classifier over the entire test set of each respective dataset and the robust accuracy ($\mathsf{RA}$) on $n=100$ randomly selected images from the test set that are correctly classified by the undefended classifier.
Our experiments have been conducted on a server equipped with two AMD EPYC 7542 CPUs, two Nvidia A40 GPUs, and 256GB RAM. The system runs Ubuntu 22.04., Python 3.9, PyTorch 1.13.0, and CUDA 11.7.
@@
(d) SurFree, ImageNet, $\mathsf{D}$=JPEG
(e) PSJA, CIFAR-100, $\mathsf{D}$=RND
(f) PSJA, CIFAR-100, $\mathsf{D}$=RCR
(g) SurFree,CIFAR-100,$\mathsf{D}$=JPEG
(h) PSJA, ImageNet, $\mathsf{D}$=RND
Figure 3: Pareto frontier when applying our approach in conjunction with RND, RCR, and JPEG on CIFAR-10, CIFAR-100, and ImageNet using the attacks PSJA and SurFree.
@@
In Figure 3, we use the Pareto frontier to showcase the tradeoff improvements of our proposal compared to the baseline relying only on the defense parameter $\nu$. We plot the points $(\mathsf{CA}(\omega^{*}),\mathsf{RA}(\omega^{*}))$ for empirically-determined Pareto-optimal solutions $\omega^{*}$ for the baseline defense ($\omega^{*}=\nu^{*}$, solid line) and for our approach ($\omega^{*}=(\nu^{*},\tau^{*})$, dotted line). The accuracy-robustness tradeoffs achieved by our proposal for selected threshold values are highlighted in grey. The specific combinations of $\nu,\tau$ that outperform the baseline tradeoffs (${\tau=1.0}$) are highlighted in Table 1 and Table 2. Note that for RCR the crop size for CIFAR-10/100 and ImageNet are shown due to the differently sized input dimensions.
@@
When evaluated against PSJA on the CIFAR-10 dataset, our proposal combined with the RND defense can provide an improvement in $\mathsf{RA}$ while preserving the  $\mathsf{CA}$ of the baseline. For example, it improves $\mathsf{RA}$ by $8\%$ (out of a maximum of $9\%$) for $\mathsf{CA}=0.95$ and $\mathsf{RA}$ by up to $9\%$ for a tradeoff of at most $1\%$ decrease in $\mathsf{CA}$. Compared to the baseline ($\tau=1.0,\nu=0.07$), we improve $\mathsf{CA}$ by $29\%$, while decreasing $\mathsf{RA}$ by just $1\%$ (cf. Figure 2(a)).
@@
Our results on CIFAR-100 shown in Figures 2(e) and 2(f) are consistent with our CIFAR-10 results. More precisely, we measure an improvement of $\mathsf{RA}$ of up to $8\%$ for a $\mathsf{CA}$ of $58\%$ when evaluating with the RND defense. Compared to the baseline with a $\mathsf{CA}/\mathsf{RA}$ of $37\%$, our approach improves the $\mathsf{CA}$ by $16\%$ and $\mathsf{RA}$ by $1\%$. When considering the RCR defense in Figure 2(f), our approach can improve $\mathsf{CA}$ by $19\%$ for the same $\mathsf{RA}$ of $43\%$. When $\mathsf{CA}=0.53$, our approach yields an improvement in $\mathsf{RA}$ by $8\%$. For ImageNet, we see an improvement of up to $1\%$ in $\mathsf{RA}$ (cf. Figure 2(h))—we discuss these results in more detail in the Appendix.
@@
Figure 2(c), Figure 2(g), and Figure 2(d) depict the evaluation results against the SurFree attack, for which our proposal can obtain consistent improvements compared to the baseline. At no decrease in $\mathsf{CA}$, we observe significant improvements in $\mathsf{RA}$. For CIFAR-10, it increases by $34\%$ and for CIFAR-100 by $21\%$. This trend is also observable for ImageNet—we obtain an increase of $\mathsf{RA}$ of $21\%$ with a negligible decrease of $\mathsf{CA}$ by $1\%$. When fixing $\mathsf{RA}$ to $67\%$, we obtain an improvement in $\mathsf{CA}$ of $4\%$ for CIFAR-10. Moreover, we see an improvement of $\mathsf{CA}$ by $3\%$ at a fixed $\mathsf{RA}=0.82$ for CIFAR-100.
@@
We now compare our proposal to various training-time defenses, which require the costly retraining of a model and cannot be retrofitted to existing models. We compare against OAT (Wang et al. 2020) because it is the only reconfigurable, i.e., adjustment of accuracy-robustness tradeoff after training, adversarial-training solution with an open-source implementation. In addition, we include traditional, non-adjustable, training-time defenses, such as PNI (He, Rakin, and Fan 2019), RSE (Liu et al. 2018), and a state-of-the-art robust model (AT) (Gowal et al. 2020) from RobustBench for further comparison. We also include results from the previously considered vanilla inference-time defenses RCR and JPEG with $\tau=1.0$, i.e., traditional deployment of defenses for this comparison.
Training-time Inference-time
$\sim$ Baseline
@@
Table 3: $\mathsf{RA}$ for state-of-the-art defenses on CIFAR-10. The baseline instantiates PSJA with $\mathsf{D}$ = RCR and SurFree with $\mathsf{D}$ = JPEG.
In Table 3, we evaluate $\mathsf{RA}$ at those values of $\mathsf{CA}$ obtained with these defenses against PSJA and SurFree on CIFAR-10. Note that some values of $\mathsf{CA}$ cannot be achieved with some defenses. In these cases, we report the $\mathsf{RA}$ that results in the closest (by up to $1-2\%$) $\mathsf{CA}$; when such a close estimate cannot be found, we do not report any value for $\mathsf{RA}$.
Against SurFree, our proposal outperforms all training-/inference-time-based defenses, outperforming AT by $8\%$, while being completely training-free. For $\mathsf{CA}=0.88$, we observe a difference to the $\mathsf{RA}$ of OAT of just $7\%$. For ${\mathsf{CA}=0.84}$, a larger difference of $23\%$ in $\mathsf{RA}$ to the one of PNI can be measured. We conclude that, although it is training-free, our proposal manages to outperform or closely perform similarly to existing defenses against SurFree. In the case of PSJA, training-time defenses outperform our approach by a minimum of $31\%$ in $\mathsf{RA}$. Compared to OAT, this difference is reduced to $24\%$. Our approach, however, results in significant improvements compared to all training-free defenses.
@@
Our proposal relies on two main parameters: the confidence threshold $\tau$ and the genuine defense’s parameter $\nu$. To evaluate the impact of each of these parameters on robustness, we present an ablation study. We detail our results in Table 1 and Table 2, where we show the achieved $\mathsf{CA}$ (top of each row) and $\mathsf{RA}$ (bottom of each row) for different values of $\tau$ and $\nu$. We highlight the values that result in an improvement of the Pareto frontier in bold. When $\tau=1.0$, our proposal instantiates the baseline since the defense is always enabled.
@@
We vary the value of $\tau\in\{0.0,0.3,0.4,0.5,0.6,0.7,0.8,0.9,0.97,0.99,1.0\}$. For the CIFAR-10 dataset, we first consider the RND defense with noise $\nu=0.05$, which offers the highest $\mathsf{RA}=0.71$ at $\mathsf{CA}=0.82$ and the best tradeoffs. As we reduce the value of $\tau$, we observe a significant improvement in $\mathsf{CA}$, increasing from $82\%$ when $\tau\geq 0.99$ to $87\%$ when $\tau=0.97$. By further reducing $\tau$ to $0.8$, our proposal achieves $94\%$ $\mathsf{CA}$, fairly close to the original $\mathsf{CA}$ of $95\%$. On the other hand, $\mathsf{RA}$ starts at $71\%$ for the baseline, remains consistent at $71\%$ when $\tau=0.99$, and decreases to $64\%$ when $\tau=0.97$. For $\tau\leq 0.8$, $\mathsf{RA}$ remains at $59\%$ and then drops sharply to a $\mathsf{RA}$ as low as $43\%$. The Pareto frontier for the fixed $\nu$ and varying $\tau$ can be found in Figure 3(a).
For RCR, we show the Pareto frontier for $\nu=22$ in Figure 3(c) and for JPEG for $\nu=85$ in Figure 3(e). For both defenses, we observe consistent improvements in the baseline tradeoffs. We argue that such a noise level introduced by these two defenses is similar to $\nu=0.05$ for the RND defense. These observations are not exclusive to CIFAR-10 but can also be clearly seen with CIFAR-100 with $\nu=26$ for RCR, while it is observable for $\nu=0.02$ and $\nu=50$ for RND and JPEG, respectively.
@@
(d) PSJA, $\mathsf{D}$=RCR, $\tau=0.7$
(e) SurFree, $\mathsf{D}$=JPEG, $\nu=85$
@@
Figure 4: Pareto frontier for various $\tau$ and $\nu$ for CIFAR-10.